A hamburger producing company’s variable cost per hamburger Is 15.50 taka and fixed cost is
2, 20,000/- taka. The company’s expected sales of hamburger are 40,000 pieces. Based on this context,
answer the following questions:

Q1: How many hamburgers the company has to sell to achieve neither profit nor loss situation? (show
the step by step calculation)

Q2: What is the minimum number of hamburgers the company has to sell to make profit?

Q3: How many hamburgers the company has to sell to incur minimum amount of loss?
1
Expert's answer
2020-09-01T11:10:45-0400

Q1

Based on the data presented in the condition, namely, the average variable costs per unit of production in the amount of 15.50 and the predicted number of products 40,000, we find the variable costs


"15.50\times 40000=620000"

We put the total costs as the sum of variable and fixed costs


"620000+220000=840000"

Determine the minimum price of products at which the company will receive neither profit nor loss

"\frac {840000}{40000}=21""MC=p"

"TC=\int MC=21Q"

"21Q=840000"


"Q=40000"

Q2:

More than 40,000


Q3:


"\frac {220000}{15.5}=14194"